OpenStudy (anonymous):

Determine whether the two graphs of the two equations are perpendicular x+2y=5 2x+4y=9

OpenStudy (anonymous):

You need to figure out the slope of each equation. You can do this by putting them into slope-intercept form y = mx + b Then compare them. If they are negative reciprocals of each other, then they are perpendicular.

OpenStudy (anonymous):

id say no

OpenStudy (anonymous):

I put them into a graph, and I still do not know. The lines are very close together

OpenStudy (anonymous):

They are perpendicular if mm' = -1.

OpenStudy (anonymous):

Can you tell me if I am right by saying no, they are not perpendicular

OpenStudy (anonymous):

Although I'm not a 100% sure, I'd say it's no.

OpenStudy (anonymous):

I agree, they are parallel. When solving for (x) you will find both have matching slopes of -1/2. Therefore they will never be perpendicular to each other.

OpenStudy (anonymous):

x + 2y = 5 2y = 5 - x y = 5/2 - x/2 Here, slope is equal to -1/2 2x + 4y = 9 4y = 9 - 2x y = 9/4 - x/2 Here, slope is equal to -1/2 Since the slopes are equal the lines are parallel. For them to be perpendicular, one would have to have a slope of 2.